# MultiScale Contextual Bandits for Long Term Objectives

### v1

@@ Appendix C Extending MSBL to $k$ levels
0:  $k\geq 2$
1:  Procedure PolicyLearning($\pi^{L(k)}_{0},\pi^{L(k-1)}_{0}$
2:     if $k=1$ then
3:        return {Base case: when at lowest level}
4:     end if
5:     PolicyLearning($\pi^{L(k-1)}_{0},\pi^{L(k-2)}_{0}$)
6:     Collect Micro Logged dataset $D^{L(k-1)}:=\{(x^{L(k-1)}_{i},a^{L(k-1)}_{i},r^{L(k-1)}_{i},p^{{}^{L(k-1)}}_{i})\}_{i=1}^{n^{{}^{L(k-1)}}}\sim\pi^{{}^{L(k-1)}}_{0}$
7:     Learn Micro policies $\hat{\Pi}^{L(k-1)}$(Eq. (5) or (6) using $D^{L1}$)
8:     Collect Macro Logged dataset $D^{L(k)}:=\{(x^{L(k)}_{j},a^{L(k)}_{j},r^{L(k)}_{j},p^{L(k)}_{j})\}_{j=1}^{n^{L(k)}}\sim\pi^{L(k)}_{0}$
9:     Learn Macro Policy $\hat{\pi}^{L(k)}\leftarrow\operatorname*{\arg\max}_{\pi^{L(k)}}\hat{V}^{L(k)}(\pi^{L(k)}\mid\hat{\Pi}^{L(k-1)};D^{L(k)})$ (Eq. (3))
10:     return learnt policies $\hat{\pi}^{L(k)},\hat{\Pi}^{L(k-1)}$
11:  end Procedure
Appendix D Experiment Details
For training the bandit policy at any level, we use Importance Sampling estimator (IPS) (Swaminathan & Joachims, 2015).
@@ Appendix D Experiment Details
For feedback modification, we additionally pass the macro action with the features, so that for Level 1 policy training with feedback modification, we have
@@
D.1 Recommender System
For training the lower level policy, which is a transformer model that uses input as the item and user embedding and predicts ranking scores, we use the user features and pre-processing following (Zhao et al., 2023).
For the upper level bandit policy, we use 5 user features, namely ‘uf_user_active_degree’,‘uf_is_live_streamer’,‘uf_is_video_author’, ‘uf_follow_user_num_range’, ‘uf_fans_user_num_range’, with their definitions provided in (Gao et al., 2022).
Increasing Level 1 and Level 2 action space. Figure 8 shows the tradeoff with varying groups in $\{2,3,4\}$ and with varying Level 1 actions as the ranking size top-$k$ varies for $k\in\{5,10,20,40\}$. We observe that the return rate for MSBL remains high across all ranking sizes and for all groups. Since we report average clicks per user, there is a decrease in the short term (clicks) with increasing ranking size. This experiment demonstrates that even with large micro action space $\mathcal{A}^{L1}$, MSBL can effectively leverage the interventions to drive the system toward the long-term reward.
(a) 2 groups
(b) 3 groups
(c) 4 groups
Figure 8: Recommender system: Tradeoff between long term return rate and clicks by varying ranking size using the boost $a^{L2}$ to the ranking as policy modification across 5 random seeds when the number of groups vary in (a) 2 groups (b) 3 groups and (c) 4 groups .
D.2 Text generation simulator
Experiment Setup
At each of the three levels, we simulate 5-dimensional contexts sampled from a normal distribution $\mathcal{N}(\mu_{g}^{l},\sigma_{f})$ belonging to two user groups. The mean vector $\mu_{g}^{l}$ is fixed according to the user group membership and the group membership is only used for the corresponding reward simulation. A default value of $\sigma_{f}=0.1$ is used. The context $x^{L1}$ represents users’ demographics and cuisine preferences. At Level 1, there are 10 bandit actions $a^{L1}$, corresponding to a cuisine:{“Ethiopian”, “Mexican”, ”French”, “Japanese”, “Spicy Indian”, “Thai”, “Carribean”, “Peruvian”, “Russian”, “Italian”, }. We use Llama-2-7b-chat (Touvron et al., 2023) as the LLM policy $\pi^{LLM}$ that acts as a personalized agent, generating response suggestions $y$ for cuisines. Each query consists of a system prompt specifying the agent’s cuisine expertise (e.g., french food, ethopian etc.) and a question $q$ asking for a cuisine suggestion. A bandit policy $\pi^{L1}(a^{L1}|x^{L1})$ selects the LLM agent system prompt $a^{L1}$. Next, we generate response $y_{t}\sim\pi^{\text{LLM}}(.|(a^{L1}\cup q)\cup y_{t-1})$ from the LLM policy, given prompt selection $a^{L1}$ and appending the previous timestep response $y_{t-1}$ to the current query. The reward is simulated as inverse perplexity conditional on the optimal prompt, representing the engagement (relevance) metric, as follows
$r^{L1}_{t}=\text{perplexity}(y_{t}|a^{L1*})^{-1}=\exp\bigg{(}\frac{1}{\text{No. of tokens in response }y_{t}}\sum_{i}^{\text{No. of tokens in response}~{}y_{t}}\log\text{prob}(y_{t,i}|y_{t,0:i-1},a^{L1*})\bigg{)}$
The query $q$ at every timestep is as follows,
You are my personal chef experienced in {cuisine} Cuisine. Your responses should be professional and concise (100 words or less). What should I eat today?
At the second level, we simulate two user groups with context $x^{L2}$ representing preferences for engagement (relevance) and diversity. We use the notion of $n$-gram repeats over multiple responses ($t1\in\{1,\ldots 10\}$) for computing diversity, where we use $n=3$. In particular, we define diversity as 1-Normalized $3$-gram repeats, where the $3$-gram repeats are computed over $T=10$ responses generated at Level 1 for each user. We simulate the reward
@@ D.2 Text generation simulator
At the third level, we simulate users with context $x^{L3}$, with one of the groups having a long-term goal of not being addicted to the app by the optimal weekly return rate. We use feedback modification during Level 2 training to get a family of policies $\hat{\Pi}^{L2}:=\hat{\pi}^{L2}_{a^{L3}}:a^{L3}\in\{[0,1],[1,0]\}\}$. The parameterized feedback is given by $\sum_{i=1}^{2}a^{L3}_{i}R_{i}$, for two reward models $R_{1},R_{2}$, such that $R_{1}=r^{L2}$, and $R_{2}=\min(\tau,r^{L2})$. We use $\tau=0.8$ as a user group-specific threshold. This simulates the scenario where for one user group, optimizing the weekly return rate is beneficial, while for the other user group, optimizing only upto a threshold $\tau$ is beneficial. We simulate $r^{L3}$ across 2 timescales of level 2 as follows,
[…]
For the second term of $r^{L3}$, we simulate a non-linear function of the weekly return rate. Specifically, activity preference is simulated as
activity $\text{activity preference}=\begin{cases}1.0&\text{if }\text{count}=1\\ 0.5&\text{if }\text{count}=2\\ 0&\text{otherwise}\end{cases}\text{, where count}=\sum_{t2}\mathbb{I}(r^{L2}_{t2}\geq\tau)$
We use $\gamma_{u}\in\{1,0\}$ as the user’s weekly return rate-long term activity preference trade-off parameter. This simulates the scenario where for one user group, the long term reward $r^{L3}$ is aligned with medium term cumulative reward $r^{L2}$. For the other group, the long and medium term are not aligned and their preference is to visit the app only once every two weeks.
Figure 9: Text generation simulator: Tradeoff between longer-term Level 2 and short-term Level 1 rewards using decoding temperature $a^{L2}\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$ as policy modification across 5 random seeds. (a) for all users (b) for users belonging to group A that prefer relevance (c) for users belonging to group B that prefer diversity among multiple responses
Figure 10: Text generation simulator: Tradeoff in different timescale rewards for the two user groups at Level 3. (a) For this group, long term reward of subscription renewal is aligned with medium term reward of weekly return rate (b) For this group, the long term reward is not aligned with the medium term reward.
Figure 11: Text generation simulator: MultiScale contexts ($x^{L2}\mid x^{L3}$) and rewards (a) group-1 where $x^{L2}$ prefers relevance given $x^{L3}$ has long term and medium term preferences aligned (b) group-2 where $x^{L2}$ prefers diversity given $x^{L3}$ has long term and medium term preferences aligned (c) group-3 where $x^{L2}$ prefers relevance given $x^{L3}$ has long term and medium term not aligned (d) group-4 where $x^{L2}$ prefers diversity given $x^{L3}$ has long term and medium term not aligned.
MultiScale contexts. In this experiment, we evaluate across contexts and timescales. At each level, we simulate contexts belonging to two user groups. Figure 9 shows the tradeoff in longer term level 2 reward and shorter term reward of clicks for the two groups that $x^{L2}$ belongs to at Level 2. Figure 10 shows the tradeoff in different timescale rewards for the two groups that $x^{L3}$ belongs to at Level 3. In both cases, we find that MSBL policy achieves near-optimal long term rewards for each of the groups. At Level 2, a context $x^{L2}$ is assigned to a user, given Level 3 context $x^{L3}$. For example, given that a user does not want to be on the app every week (Level 3 preference), the user likes more diverse responses (Level 2 preference). Figure 11 shows the tradeoff in expected rewards across 4 groups formed by contexts at Level 2 and 3. While there are user groups for whom a policy that is optimal in the medium term is also near-optimal in the long term (groups 1 and 3 with overlapping red and blue lines), Level 3 opt. (MSBL) pareto dominates each of these policies for all 4 groups for the long term.
D.3 Multi-turn conversation
[…]
@@ D.3 Multi-turn conversation
Keep your response focused and relevant to the ongoing discussion.
Do not include any meta-text like User:’, Human:’, or explanatory phrases. Your entire response should be just the message itself.
Do not generate the assistants response - only provide the user(Human)’s next message.

### v2

@@ D.2 Extending MSBL to $k$ levels
Appendix E Multi-Scale Policy Framework
The PAC Bayes generalization bounds we consider in Eq. (3) are derived for counterfactual risk minimization (e.g., clipped inverse propensity estimator) in [17].
As part of the motivation, we illustrate the following example.
E.1 Gaussian Model
We consider Gaussian parameterizations, where each policy $\pi(.|.,\theta)\in\Pi$ is defined via a parameter vector $\theta\in\operatorname{\mathbb{R}}^{d}$. We define the target policy distribution as $Q^{L2*}=N(\theta^{L2*},\Sigma^{L2})$, an uninformed prior distribution $P_{0}=N(\theta_{0},\Sigma_{0})$ for some arbitrary $\theta_{0}$, and an informed prior $P^{L1}=N(\theta^{L1},\Sigma^{L1})$ centered at the learned micro policy $\hat{\pi}^{L1}$.
The KL divergence between any two multivariate gaussian $KL(Q^{L2*}||P_{0})$ is given by,
$KL(Q^{L2*}||P_{0})=\frac{1}{2}\left[\log\frac{|\Sigma_{0}|}{|\Sigma^{L2}|}-d+\text{tr}(\Sigma_{0}^{-1}\Sigma^{L2})+(\theta_{0}-\theta^{L2})^{T}\Sigma_{0}^{-1}(\theta_{0}-\theta^{L2})\right]$
where $(\theta_{0}-\theta^{L2})^{T}(\Sigma_{0})^{-1}(\theta_{0}-\theta^{L2})=|\theta_{0}-\theta^{L2}|_{M}$ is the squared Mahalanobis distance in the parameter space $\theta$.
Using the above, the gain in the number of samples from using an informed prior $P^{L1}$ instead of an uninformed prior $P_{0}$ is,
$n_{0}-n^{L2}\propto KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})=$
$\text{tr}(\Sigma_{0}^{-1}\Sigma^{L2})-\text{tr}({{(\Sigma^{L1})}^{-1}}\Sigma^{L2})+(\theta^{L2}-\theta_{0})^{T}\Sigma_{0}^{-1}(\theta^{L2}-\theta_{0})-(\theta^{L2}-\theta^{L1})^{T}{{(\Sigma^{L1})}^{-1}}(\theta^{L2}-\theta^{L1})$
$+\log\frac{|\Sigma_{0}|}{|{\Sigma}^{L1}|}$ (8)
Setting $\Sigma^{L1}:=\Sigma_{0}=\Sigma_{P}$ and since the squared Mahalanobis distance is symmetric, we have
$KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})=(\theta^{L2}-\theta_{0})^{T}\Sigma_{P}^{-1}(\theta^{L2}-\theta_{0})-(\theta^{L2}-\theta^{L1})^{T}{\Sigma_{P}^{-1}}(\theta^{L2}-\theta^{L1})$
As a result, we can have the sample gain from using the informed prior $P^{L1}$ instead of $P_{0}$ as
$n_{0}-n^{L2}\propto KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})\in O(|\theta^{L2*}-\theta_{0}|_{M}-|\theta^{L2*}-\theta^{L1}|_{M})$ (9)
For the special case of isotropic Gaussian distributions, where $\Sigma^{L1}=\sigma^{L1}\mathbb{I}:=\sigma_{0}\mathbb{I}=\sigma_{P}\mathbb{I}$ , Eq. (9) can also be written in terms of L2 distance as,
$KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})\in O(||\theta^{L2*}-\theta_{0}||^{2}-||\theta^{L2*}-\theta^{L1}||^{2})$
However, the above is a conservative estimate, since it can be beneficial to pick variances $\sigma^{L1}<\sigma_{0}$ that increasingly reduce the variance going from uninformed prior $P_{0}$ to $P^{L1}$.
Next, we simulate a toy example with isotropic Gaussian distributions for $\theta\in\operatorname{\mathbb{R}}^{50}$. To calculate $n_{0}-n^{L1}$, we use the exact KL divergence in Eq. (8). We start with an uninformed prior variance $\sigma_{0}=200$ and learn all 50 parameters with target variance $\sigma^{L2}=1.0$. For $n^{L2}$, we start with an informed prior $P^{L1}$ that has 49 parameters learned with $\sigma^{L1}=1.0$, and we only need to adjust 1 more parameter. The coefficient constant for $\frac{n}{KL(Q||P)}$ is used as $5.0e^{3}$ for all cases. This provides $\frac{n_{0}-n^{L2}}{n_{0}}=98\%$ as the reduction in samples needed at the macro level L2. Note that the number of samples used to form the prior $n^{L1}$ is significantly cheaper than $n^{L2}$ since they occur $T$ times as frequently. Empirically, even when taking the additional L1 samples into account, this still provides a $\frac{n_{0}-n^{L2}-\frac{n^{L1}}{T}}{n_{0}}=88.2\%$ reduction with a $T=10$ horizon.
Appendix F Experiment Details
For training the bandit policy at a given level, we use Importance Sampling estimator (IPS) [33].
@@ Appendix F Experiment Details
For feedback modification, we additionally pass the macro action with the features, so that,
@@
Figure 5: MultiScale Simulator Modules
Code and Simulator The simulator code to reproduce experimental results can be found at https://github.com/RichRast/mspl. Figure 5 shows the structure of the simulator with four key modules. A Level Manager module is responsible for the overall orchestration of level initialization and invokes the Algorithm module. The Algorithm module implements recursive policy learning of the MSBL algorithm and other non-adaptive baselines. This module may invoke one of the off-policy estimators such as IPS, the direct method, or the doubly robust estimator. Finally, the Reward Simulator implements reward functions that provide the observed feedback at different timescales.
Policy learning options for the algorithm, estimators, and the type of reward function are configurable among the existing options, via the user config file. Further, each of the Algorithm, Estimators, and Reward Simulator modules can be customized to add new functionality.
Compute Resources We use NVIDIA RTX A6000-48GB GPUs for experiments in conversational systems and 24GB NVIDIA GeForce RTX 3090 for experiments in conventional recommender systems.
F.1 A toy example to illustrate comparison with flat RL
We consider a finite horizon setting MDP with the goal of learning $\pi(a|x)$ that optimizes the long term reward $r^{L2}\sim p(r^{L2}|(x_{i},a_{i}),\ldots(x_{i+T},a_{i+T}))$ observed after $T$ timesteps.
Figure 6: MSBL episodic samples vs Q-learning episodic samples for a toy setup as the action space increases with $k$.
We use (tabular) Q learning to learn a policy that optimizes for the sparse reward $r^{L2}$ for the following setup. Users arrive with context $x_{t}\sim p(x_{t})$ and select $k$ out of $n$ items based on relevance vector $r_{t}(x_{t})\in\operatorname{\mathbb{R}}^{n}$. The action $a_{t}$ can be one out of $k$ choose $n$ combinations. We simulate $n=10$ items, and 2 contexts $x_{t}$. The long term reward for each user is observed based on their preference for the item. We use a time horizon of $T=5$.
While Q learning is provably optimal, it is computationally expensive to form the Q table by learning from episodes $\{(x_{i},a_{i}),\ldots(x_{i+T},a_{i+T}),r^{L2}\}$, especially when $r^{L2}$ is sparse. In contrast, MSBL leverages the problem structure as follows. Its micro policy is simply argmax on item relevances. Even though this myopic policy doesn’t lead to optimal long-term reward, it provides a prior for the macro level, which only needs to learn the macro intervention $a^{L2}$ as the amount of boost to the relevance vector $r_{t}$. We construct $|\mathcal{A}^{L2}|=8$, so that $\pi^{L2}$ only needs to select the best $a^{L2}$ for each user context.
As a side note, while function approximation and other advancements (actor-critic, etc.) would provide benefit to the RL baseline, they are orthogonal to our key idea of enabling faster learning via the use of hierarchical priors for sparse long-term feedback. Figure 6 shows the number of RL episodic samples $n_{0}$ on the x-axis. On the y-axis, we plot the MSBL macro policy samples $n^{L2}$ to achieve the same $r^{L2}$ in expectation as that of Q-learning. We can see that $10\leq\frac{n_{0}}{n^{L2}}\leq 62$ as $k$ varies in $\{2,4,6\}$. Scaling this baseline for continuous context features, time horizon, $k$, and number of items $n$ is challenging and motivates our approach.
F.2 Multi-turn conversation
@@
Figure 7: Long-term (multi-turn) vs short-term (single-turn) rewards for all users (overall) and for each user group.
We use the upper level reward simulator as Llama-3-70b API [12], which scores user satisfaction for the given user persona and full 5-turn conversation as a binary score in $\{0,1\}$. We transform the user persona text with MPNet model [31] and project down with PCA to a 10-dimensional embedding. This forms the upper-level context $x^{L2}$ for bandit policy $\pi^{L2}$.
Group-wise Results Figure 7 shows the long term multi turn reward vs per-conversation short term reward for each of the user personas - Child and Expert. We note that MSBL achieves the best multi-turn reward for each user persona while the other policies are optimal only for one type of user group.
@@
F.3 Conversational recommender system
Experiment Setup At each of the three levels, we simulate 5-dimensional contexts sampled from a normal distribution $\mathcal{N}(\mu_{g}^{l},\sigma_{f})$ belonging to two user groups. The context $x^{L1}$ represents users’ demographics and cuisine preferences. At Level 1, there are 10 bandit actions $a^{L1}$, corresponding to a cuisine:{“Ethiopian”, “Mexican”, ”French”, “Japanese”, “Spicy Indian”, “Thai”, “Carribean”, “Peruvian”, “Russian”, “Italian”, }. We use Llama-2-7b-chat [34] as the LLM policy $\pi^{LLM}$ that acts as a personalized agent, generating cuisine suggestions $y$ to users at L1 timescale $t1=\{1,\ldots 10\}$. Each query consists of a system prompt specifying the agent’s expertise and a user query $q$. We learn a bandit policy $\pi^{L1}(a^{L1}|x^{L1})$ that selects the particular LLM agent $a^{L1}$. Next, we generate response $y_{t}\sim\pi^{\text{LLM}}(.|(a^{L1}\cup q)\cup y_{t-1})$ from the LLM policy, given the agent selection $a^{L1}$ and append the previous timestep response $y_{t-1}$ to the current query. The L1 reward is simulated as inverse perplexity conditional on the optimal prompt, representing the engagement (relevance) metric, as follows
$r^{L1}_{t}=\text{perplexity}(y_{t}|a^{L1*})^{-1}=\exp\bigg{(}\frac{1}{\text{\#tokens in response }y_{t}}\sum_{i}^{\text{\# tokens in}~{}y_{t}}\log\text{prob}(y_{t,i}|y_{t,0:i-1},a^{L1*})\bigg{)}$
The query $q$ at every timestep $t\geq 1$ is as follows,
You are my personal chef experienced in {cuisine} Cuisine. Your responses should be professional and concise (100 words or less). Previously, you suggested: {previous generated response}.
What should I eat today?
At the second level, we simulate two user groups with context $x^{L2}$ representing preferences for engagement (relevance) and diversity. We use the notion of $n$-gram repeats over multiple responses ($t1\in\{1,\ldots 10\}$) for computing diversity, where we use $n=3$. In particular, we define diversity as normalized $3$-gram repeats, where the $3$-gram repeats are computed over $T=10$ responses generated at Level 1 for each user. We simulate the reward
[…]
Some sample responses (decoding temperatures are logged after generation) are,
"Hola! How about our delicious carne asada burritos with sauted onions and bell peppers, topped with melted cheese and served with a side of sliced avocado and a drizzle of homemade salsa? decoding_temp_1.10",
"For lunch today, I recommend our Chili con Carne Burrito.’ Its a satisfying blend of slow-cooked beef, beans, and vegetables wrapped in a large flour tortilla, topped with creamy sour cream and fresh cilantro. Its a hearty and flavorful dish that will warm you up on a chilly day! " decoding_temp_0.87",
"Great! Based on your previous preferences, I recommend trying our Shiro Wot - a creamy stew made with chickpeas, onions, garlic, and a blend of spices. Its a popular vegetarian dish thats sure to satisfy your taste buds. Pair it with a side of Kik Alicha - sauted vegetables in a mild sauce, and enjoy! decoding_temp_0.21",
At the third level, we simulate users with context $x^{L3}$. We use feedback modification for Level 2 training to get a family of policies $\hat{\Pi}^{L2}:=\{\hat{\pi}^{L2}_{a^{L3}}:a^{L3}\in\{[0,1],[1,0]\}\}$. The parameterized feedback is given by $\sum_{i=1}^{2}a^{L3}_{i}R_{i}$, for two reward models $R_{1},R_{2}$, such that $R_{1}=r^{L2}$, and $R_{2}=\min(\tau,r^{L2})$. We use $\tau=0.8$ as a user group-specific threshold. This simulates the scenario where for one user group, optimizing the weekly return rate is beneficial, while for the other user group, optimizing only upto a threshold $\tau$ is beneficial. We simulate $r^{L3}$ across 2 timescales of level 2 as follows,
[…]
where $\gamma_{u}\in\{1,0\}$ is the trade-off parameter.
Figure 9: Tradeoff in different timescale rewards for the two user groups at Level 3. (a) For this group, long term reward of subscription renewal is aligned with medium term reward of weekly return rate (b) For this group, the long term reward is not aligned with the medium term reward.
Figure 10: MultiScale contexts ($x^{L2}\mid x^{L3}$) and rewards for 4 groups in (a), (b), (c), and (d) across all three timescales.
MultiScale contexts. In this experiment, we evaluate across user groups from all three timescales. At each level, we simulate contexts belonging to two user groups.
Figure LABEL:fig:LLM_2levels_groupwise shows the tradeoff in longer-term L2 reward and shorter term L1 reward for the two user groups individually. The context $x^{L2}$ at Level 2 belongs to these two groups. Similarly, Figure 9 shows the tradeoff in different timescale rewards for the two groups that $x^{L3}$ belongs to at Level 3. In both cases, we find that MSBL policy achieves near-optimal long term rewards for each of the groups.
A Level 2 context $x^{L2}$ is drawn given a Level 3 context $x^{L3}$. For example, given that a user does not want to be on the app every week (Level 3 preference), the user likes more diverse responses (Level 2 preference). This results in 4 user groups across L2 and L3. Figure 10 shows the tradeoff in expected rewards across these 4 groups. While there are user groups for whom a policy that is optimal in the medium term is also near-optimal in the long term (groups 1 and 2 with overlapping red and blue lines), Level 3 opt. (MSBL) pareto dominates each of these policies for all the 4 groups for the long term reward.
Policy
Noise MSBL $a^{L2}\sim\pi_{0}^{L2}$ $a^{L2}_{1}$ $a^{L2}_{2}$ $a^{L2}_{3}$ $a^{L2}_{4}$ $a^{L2}_{5}$ $a^{L2}_{6}$
0.05 0.82$\pm$0.13 0.66$\pm$0.10 0.46$\pm$0.13 0.68$\pm$0.16 0.76$\pm$0.11 0.77$\pm$0.10 0.72$\pm$0.08 0.64$\pm$0.06
0.10 0.86$\pm$0.08 0.66$\pm$0.09 0.48$\pm$0.10 0.61$\pm$0.13 0.77$\pm$0.12 0.74$\pm$0.11 0.75$\pm$0.11 0.62$\pm$0.06
0.20 0.76$\pm$0.10 0.57$\pm$0.08 0.43$\pm$0.18 0.59$\pm$0.10 0.60$\pm$0.08 0.61$\pm$0.11 0.58$\pm$0.08 0.55$\pm$0.06
0.30 0.66$\pm$0.12 0.54$\pm$0.06 0.34$\pm$0.17 0.49$\pm$0.12 0.49$\pm$0.12 0.55$\pm$0.03 0.53$\pm$0.04 0.53$\pm$0.02
0.40 0.58$\pm$0.06 0.53$\pm$0.02 0.37$\pm$0.09 0.45$\pm$0.12 0.51$\pm$0.02 0.51$\pm$0.04 0.52$\pm$0.01 0.51$\pm$0.01
0.50 0.62$\pm$0.08 0.47$\pm$0.03 0.33$\pm$0.09 0.50$\pm$0.05 0.51$\pm$0.01 0.52$\pm$0.02 0.52$\pm$0.01 0.51$\pm$0.01
Table 1: Level 2 Reward with varying noise in features
Robustness to noisy features. We also evaluate robustness of MSBL with varying noise in the user features with $\sigma_{f}\in\{0.05,\boldsymbol{0.1},0.2,0.3,0.4,0.5\}$ with default value in bold. Table 1 shows that while the expected Level 2 rewards decrease with increasing noise in features, MSBL maintains the advantage of learning the interventions over other baselines across noise variations.
F.4 Recommender System
In the final setting, we use the KuaiRand dataset [11] to simulate two levels of short and long term feedback. This allows us to evaluate based on real-world item, user features, and their historical interactions, thereby evaluating for increasing action spaces. We use the simulator developed in [38] and modify it for the contextual bandit setting. We use a training dataset of size 7,829 and test randomly selected 1,174 users. There are 5659 items and their descriptions/features in the dataset. At the lower level, each user logs $T=5$ interactions, and a transformer model uses item and user embeddings to predict ranking scores $s$. Action $a^{L1}$ represents a top-k ranking of items for a context $x^{L1}$, from the policy $\hat{\pi}^{L1}\leftarrow\operatorname*{\arg\max}_{k}s$. Then, we simulate clicks according to a Bernoulli distribution of the ranking scores of selected items and micro-level reward as average clicks per user.
For training the lower level policy, which is a transformer model that uses input as the item and user embedding and predicts ranking scores, we use the user features and pre-processing following [38].
For the upper level bandit policy, we use 5 user features, namely ‘uf_user_active_degree’,‘uf_is_live_streamer’,‘uf_is_video_author’, ‘uf_follow_user_num_range’, ‘uf_fans_user_num_range’, with their definitions provided in [11].
Increasing Level 1 and Level 2 action space. Figure LABEL:fig:tradeoff_recSys_slate shows the tradeoff with varying groups in $\{2,3,4\}$ and with varying Level 1 actions as the ranking size top-$k$ varies for $k\in\{5,10,20,40\}$. We observe that the return rate for MSBL remains high across all ranking sizes and for all groups. Since we report average clicks per user, there is a decrease in the short term (clicks) with increasing ranking size. This experiment demonstrates that even with large micro action space $\mathcal{A}^{L1}$, MSBL can effectively leverage the interventions to drive the system toward the long-term reward.
Figure 13: Long term return rate with varying noise $\sigma_{s}$ in micro policy varies in $\{0.0,0.2,0.4,0.6,0.8,1.0,2.0\}$.
Robustness to errors in the micro policy. To evaluate the robustness with varying errors in the micro policy, we simulate a perturbed policy $\tilde{\pi}^{L1}_{a^{L2}}:=\operatorname*{\arg\max}_{k}\tilde{s}$, where noisy ranking scores $\tilde{s}$ are sampled from $\mathcal{N}(s,\sigma^{2}_{s})$, with $\sigma_{s}\in\{0.0,0.2,0.4,0.6,0.8,1.0,2.0\}$. Figure 13 shows long term reward as this noise varies for top-10 ranking with 2 groups. We find that learning the macro policy is robust for fairly large noise $\sigma_{s}$ up to $1.0$. In real-world deployment, the micro policy may be updated periodically, so this robustness to errors makes the nested bandit learning practically appealing.